OpenStudy (anonymous):

A parcel delivery service will deliver a package only if the length plus the girth (distance around, taken perpendicular to the length) does not exceed 100 inches. Find the maximum volume of a rectangular box with square ends that satisfies the delivery company's requirements.

OpenStudy (anonymous):

L +2W+2D =< 100 for a rectangle with square ends D=W L+4W =<100 V= LW^2 use L=100 -4W to sub.s in V= (100-4W)W^2 Vmax is at: V' = 0 = 200W -12W^2 solve for W use W to find L use L and W to find Vmax.